OpenStudy (anonymous):

The slope of a line is –4 and its y-intercept is (0, 4). What is the equation of the line that is perpendicular to the first line and passes through (–8, 2)?

OpenStudy (anonymous):

If we want our line was perpendicular to a line we have this : New m=-1/m1 Means : m.m1=-1 and our m is 1/4 Got it ?

OpenStudy (anonymous):

There is a law in mathematical statistics if a line is perpendicular to a line then the products of their respective gradients is -1

OpenStudy (anonymous):

therefore the gradient of the line = -4 * x = -1 x= 1/4

OpenStudy (anonymous):

y = 0.25x +c (where c is the y intercept)

OpenStudy (anonymous):

I think I said that ??!

OpenStudy (anonymous):

By substituting (–8, 2) these co- ordinates into the equation u will get the answer in the form of y = mx + C

OpenStudy (anonymous):

thank you guys i understand now